OpenStudy (anonymous):

The denominator of the fraction 4/13 is increased by an amount so that the value of the resulting fraction is 2/9. Determine the amount by which the denominator was increased.

OpenStudy (anonymous):

4/(13+x)=2/9 cross multiply 2(13+x)=9*4 can you solve from here?

OpenStudy (anonymous):

is it 5? :D